If the sum of the mean and variance of a Binomial distribution of 5 trials is 9/5,find p(x>=1).
1
Expert's answer
2019-08-29T08:34:27-0400

The mean and variance of the Binomial distribution are given by the following formulas: "\mu=np" and "\sigma^2=np(1-p)" .

5 trials means that "n=5" .

Substitute this value into two formulas:

"\mu=5p"

"\sigma^2=5p(1-p)"

Since the sum of the mean and variance is "9\/5" :

"5p+5p(1-p)=9\/5"

This is a quadratic equation that we need to solve in order to find "p"

"5p+5p-5p^2=9\/5"

"-5p^2+10p=9\/5"

Divide both sides by "-5" :

"p^2-2p=-9\/25"

Complete the square:

"p^2-2p+1=-9\/25+1"

"(p-1)^2=16\/25"

"p-1=-4\/5" or "p-1=4\/5"

"p=1\/5=0.2" or "p=9\/5=1.8"

Since the probability can't be greater than 1, the only solution is "p=0.2" .

Now, that we know the value of "p" and "n" , we can find "P(x \ge 1)" . Since it's the complement to "P(x=0)" ,

"P(x \ge 1)=1-P(x=0)"

General formula for "P(x)" for Binomial distribution is "P(x)=C(n,x)p^x(1-p)^{n-x}" , thus:


"P(x \ge 1)=1-P(x=0)=""=1-C(5,0)\cdot 0.2^0\cdot (1-0.2)^{5-0}=""=1-1\cdot 1\cdot 0.8^5=0.67232"

Answer: "P(x \ge 1)=0.67232" .